Sand is deposited at a uniform rate of 20 Kg/s and with negligible kinetic energy on to an empty conveyor belt moving horizontally at a constant speed of 10 meters per minute. Find (a) the force required to maintain the constant velocity, (b) the power required to maintain a constant velocity, and (c) the rate of change of kinetic energy of the moving sand
1
Expert's answer
2020-11-03T10:47:05-0500

From Newton's second law, Force is defined as the rate of change in momentum.


"F=\frac{dP}{dt}=\frac{d(mv)}{dt}=m\frac{dv}{dt}+v\frac{dm}{dt}\\
F=m\frac{dv}{dt}+v\frac{dm}{dt}\hspace{2cm}(1)\\
\textsf{where P, m, v and t represent momentum,mass, velocity and time respectively.}\\"


(a) Since the conveyor belt is to maintain a constant velocity, "\frac{dv}{dt}=0."

Equation (1) therefore becomes "F=v\frac{dm}{dt}.\\ \textsf{From the question,}\\
 v=\frac{10m}{min} = \frac{10m}{60s}= \frac{1}6ms^{-1}\\
\frac{dm}{dt}=20kgs^{-1}\\\hspace{2cm}\\

\textsf{The Force is,}\\
F=\frac{1}6ms^{-1} × 20kgs^{-1}=\frac{10}3kgms^{-2}\\
F=3.33N."


(b) The power is given by,

"P=F×v
P=3.33N×\frac{1}6ms^{-1}\\
P=0.56W"


(c) From work energy principle,

Change in Kinetic energy = Workdone

"\frac{d(K.E)}{dt}=\frac{Workdone}{time}\\
\frac{d(K.E)}{dt}=Power\\
\frac{d(K.E)}{dt}=0.56W"

The rate of change of kinetic energy of the moving sand is therefore "0.56W"